# Imbalance-Robust and Sampling-Efficient Continuous Conditional GANs via Adaptive Vicinity and Auxiliary Regularization

### v1

@@ 2.1 Continuous Conditional Generative Models
where $D$ and $G$ denote the discriminator and generator networks, superscripts $r$/$g$ indicate real/fake samples, $\bm{z}_{i}$ represents Gaussian noise, $\sigma$ is a positive hyperparameter, and $N^{r}$/$N^{g}$ denote real/fake sample sizes. As illustrated in Figure 2(a), the hard vicinity defines a closed interval centered at $y_{c}$ with a fixed radius $\kappa$. Thus, the weights $W^{r}_{i}$ and $W^{g}_{i}$ for HVDL are defined as follows:
@@
with ${}\mathds{1}_{\{\cdot\}}$ as the indicator function and $N_{y_{c},\kappa}$ counting samples in the hard vicinity $[y_{c}-\kappa,y_{c}+\kappa]$. Ding et al. (2023b) provides a rule of thumb to compute $\kappa$ and $\sigma$ from samples. For soft vicinity, all samples are used to estimate $p_{r}(\bm{x}|y_{c})$, but each sample $\bm{x}_{i}$ is assigned with a weight $W_{i}$, with the following formulation:
@@ 3.2 Adaptive Vicinity for Imbalanced Data
Let $Q^{r}=\{(\bm{x}^{r}_{i},y^{r}_{i})\}_{i=1}^{N^{r}}$ denote a training set of $N^{r}$ real image-label pairs, where the labels consist of $M$ distinct values $Y^{u}=\{y_{(1)},...,y_{(M)}\}$ sorted in strictly increasing order ($y_{(i)}<y_{(i+1)}$ for $i=1,...,M$). For $i$-th distinct label value $y_{(i)}$, we denote by $N_{i}$ its corresponding sample count. In severely imbalanced datasets, $N_{i}$ exhibits substantial variation across different $i$, as demonstrated in Figure 1(b)-(c).
@@
We further propose a hybrid AV mechanism (Figs. 2(d) and 4) that extends the soft AV approach. This hybrid variant incorporates truncated weight decay by assigning zero weights to samples outside the interval $[y_{c}-\kappa_{y_{c}},y_{c}+\kappa_{y_{c}}]$. By combining the adaptive weighting of soft AV with the boundary constraints of hard vicinity, this approach can further enhances label consistency.
@@ 4.1 Experimental Setup
Datasets. Building upon established experimental protocols (Ding et al. 2023b; Ding, Wang, and Xu 2024; Ding et al. 2025), we conduct comprehensive evaluations across multiple benchmark datasets: RC-49 (Ding et al. 2023b) (44,051 chair images at $\times 64$ 64 resolution with rotation angles from $0.1^{\circ}$ to 89.989.9$\circ$89.9 ∘ in $0.1^{\circ}$ increments), UTKFace (Zhang, Song, and Qi 2017) (14,723 facial images across $\times 64$ 64, $\times 128$ 128, and $\times 192$ 192 resolutions with age labels from 1 to 60), and Steering Angle (Chen 2018) (12,271 driving images at $\times 64$ 64 and $\times 128$ 128 resolutions covering steering angles from $-80^{\circ}$ to $80^{\circ}$). For UTKFace and Steering Angle experiments, all available samples are used for both training and evaluation. The RC-49 experiments employ a curated balanced subset of 11,250 images (25 samples each for 450 distinct angle labels) for training while utilizing the complete collection of 44,051 images for comprehensive evaluation.
@@
Evaluation Setup. For systematic evaluation, each trained model generates a predetermined number of images at specified evaluation centers across all datasets. On RC-49 and RC-49-I, models produce 200 fake images for each of 899 rotation angles spanning $0.1^{\circ}$ to 89.989.989.9 in $0.1^{\circ}$ increments. For UTKFace, models generate 1,000 images per age from 1 to 60 years. The Steering Angle evaluation employs 2,000 uniformly distributed points across $[-80^{\circ},80^{\circ}]$, with 50 images synthesized per point. This protocol yields total generation counts of 179,800 images (RC-49 and RC-49-I), 60,000 images (UTKFace), and 100,000 images (Steering Angle) for each evaluated method.
@@ 4.2 Experimental Results
CcGAN (T-PAMI’23) 0.413 1.733 1.329 8.240 $5$.56 2.44
Dual-NDA (AAAI’24) 0.396 1.678 1.298 6.765 $5$.56 2.44
@@
Table 1: Comparison of generation quality and sampling efficiency at $\times 64$ 64 resolution. Performance is assessed using four metrics: SFID (overall quality), NIQE (visual realism), Diversity (sample diversity), and Label Score (label consistency), along with sampling speed and GPU memory consumption measured on an RTX 4090D. $\downarrow$” ($\uparrow$”) indicates lower (higher) values are preferred. The best and second-best results are highlighted in bold and underlined, respectively.
@@
CcGAN-AVAR-S 0.455 1.621 1.104 6.433 $1$.11 13.18
CcGAN-AVAR-H 0.435 1.588 1.187 6.377 $1$.11 13.18
Table 2: Comparison of generation quality and sampling efficiency at high resolutions (>64×64>64$\times$ 64> 64 × 64) across methods. The best and second-best results are highlighted in bold and underlined, respectively.
@@ 4.3 Ablation Study
  • The Effects of $N_{\text{AV}}$: We investigate the impact of the key hyperparameter $N_{\text{AV}}$ on CcGAN-AVAR’s performance (Figure 7), conducting this ablation study with auxiliary regularization disabled to isolate AV’s effect. The results demonstrate substantial performance degradation when either disabling AV entirely ($N_{\text{AV}}=0$) or employing an excessively wide vicinity (large $N_{\text{AV}}$ values), while moderate $N_{\text{AV}}$ values consistently yield acceptable results. The dashed green lines in Figure 7 correspond to the optimal values selected for our main experiments.
@@ B.2 Relation to ACGAN
CcGAN-AVAR is fundamentally distinct from ACGAN (Odena, Olah, and Shlens 2017). First, while ACGAN is designed for class-conditional generative modeling with an auxiliary classification branch, its discriminator does not take class information as input. In contrast, CcGAN-AVAR incorporates a regression branch for continuous label prediction and explicitly feeds condition labels into the discriminator via label projection (Figure 3). Second, in terms of training objectives, ACGAN employs a cross-entropy loss for classification, whereas CcGAN-AVAR adopts a regression loss. Notably, the proposed regression loss (Eq. (6)) is not a standard mean squared error but a specifically designed $\gamma$-insensitive hinge loss, which enhances robustness to label noise. Third,
@@ C.2 Hardware and Software Details
  • $\times 64$ 64 resolution: 24GB GPU memory
  • $\times 128$ 128 and $\times 192$ 192 resolutions: 48GB GPU memory (RTX 4090D-48G)
@@ Appendix D Extra Experimental Results
Example synthetic images generated by Dual-NDA and CcGAN-AVAR for the UTKFace and Steering Angle experiments (128$\times$128 resolution) are presented in Figs. S.12 and S.13.
Figure S.12: Some example fake images from Dual-NDA and CcGAN-AVAR-H in the UTKFace (128$\times$128) experiment.
Figure S.13: Some example fake images from Dual-NDA and CcGAN-AVAR-H in the Steering Angle (128$\times$128) experiment.

### v2

@@ 2.1 Continuous Conditional Generative Models
where $D$ and $G$ denote the discriminator and generator networks, superscripts rr/$g$ indicate real/fake samples, $\bm{z}_{i}$ represents Gaussian noise, $\sigma$ is a positive hyperparameter, and $N^{r}$$N^{g}$ denote real/fake sample sizes. As illustrated in Figure 2(a), the hard vicinity defines a closed interval centered at $y_{c}$ with a fixed radius $\kappa$. Thus, the weights $W^{r}_{i}$ and $W^{g}_{i}$ for HVDL are defined as follows:
@@
with $\mathds{1}_{\{\cdot\}}$ as the indicator function and $N_{y_{c},\kappa}$ counting samples in the hard vicinity $y_{c}-\kappa,y_{c}+\kappa]$. Ding et al. (2023b) provides a rule of thumb to compute $\kappa$ and $\sigma$ from samples. For soft vicinity, all samples are used to estimate $p_{r}(\bm{x}|y_{c})$, but each sample $\bm{x}_{i}$ is assigned with a weight $W_{i}$, with the following formulation:
@@ 3.2 Adaptive Vicinity for Imbalanced Data
Let $Q^{r}=\{(\bm{x}^{r}_{i},y^{r}_{i})\}_{i=1}^{N^{r}}$ denote a training set of $N^{r}$ real image-label pairs, where the labels consist of $M$ distinct values $Y^{u}=\{y_{(1)},...,y_{(M)}\}$ sorted in strictly increasing order ($y_{(i)}<y_{(i+1)}$ for i=1,,Mi=1,...,M). For $i$-th distinct label value $y_{(i)}$, we denote by $N_{i}$ its corresponding sample count. In severely imbalanced datasets, $N_{i}$ exhibits substantial variation across different $i$, as demonstrated in Figure 1(b)-(c).
@@
We further propose a hybrid AV mechanism (Figs. 2(d) and 4) that extends the soft AV approach. This hybrid variant incorporates truncated weight decay by assigning zero weights to samples outside the interval $y_{c}-\kappa_{y_{c}},y_{c}+\kappa_{y_{c}}]$. By combining the adaptive weighting of soft AV with the boundary constraints of hard vicinity, this approach can further enhances label consistency.
@@ 4.1 Experimental Setup
Datasets. Building upon established experimental protocols (Ding et al. 2023b; Ding, Wang, and Xu 2024; Ding et al. 2025), we conduct comprehensive evaluations across multiple benchmark datasets: RC-49 (Ding et al. 2023b) (44,051 chair images at $64\times 64$ resolution with rotation angles from $0.1^{\circ}$ to $89.9\circ$ in $0.1^{\circ}$ increments), UTKFace (Zhang, Song, and Qi 2017) (14,723 facial images across $64\times 64$, $128\times 128$, and $192\times 192$ resolutions with age labels from 1 to 60), and Steering Angle (Chen 2018) (12,271 driving images at $64\times 64$ and $128\times 128$ resolutions covering steering angles from $80^{\circ}$ to $80^{\circ}$). For UTKFace and Steering Angle experiments, all available samples are used for both training and evaluation. The RC-49 experiments employ a curated balanced subset of 11,250 images (25 samples each for 450 distinct angle labels) for training while utilizing the complete collection of 44,051 images for comprehensive evaluation.
@@
Evaluation Setup. For systematic evaluation, each trained model generates a predetermined number of images at specified evaluation centers across all datasets. On RC-49 and RC-49-I, models produce 200 fake images for each of 899 rotation angles spanning $0.1^{\circ}$ to $89.9$ in $0.1^{\circ}$ increments. For UTKFace, models generate 1,000 images per age from 1 to 60 years. The Steering Angle evaluation employs 2,000 uniformly distributed points across $80^{\circ},80^{\circ}]$, with 50 images synthesized per point. This protocol yields total generation counts of 179,800 images (RC-49 and RC-49-I), 60,000 images (UTKFace), and 100,000 images (Steering Angle) for each evaluated method.
@@ 4.2 Experimental Results
CcGAN (T-PAMI’23) 0.413 1.733 1.329 8.240 5555.56 2.44
Dual-NDA (AAAI’24) 0.396 1.678 1.298 6.765 5555.56 2.44
@@
Table 1: Comparison of generation quality and sampling efficiency at $64\times 64$ resolution. Performance is assessed using four metrics: SFID (overall quality), NIQE (visual realism), Diversity (sample diversity), and Label Score (label consistency), along with sampling speed and GPU memory consumption measured on an RTX 4090D. $\downarrow$” ($\uparrow$”) indicates lower (higher) values are preferred. The best and second-best results are highlighted in bold and underlined, respectively.
@@
CcGAN-AVAR-S 0.455 1.621 1.104 6.433 1111.11 13.18
CcGAN-AVAR-H 0.435 1.588 1.187 6.377 1111.11 13.18
Table 2: Comparison of generation quality and sampling efficiency at high resolutions ($64\times 64$) across methods. The best and second-best results are highlighted in bold and underlined, respectively.
@@ 4.3 Ablation Study
  • The Effects of $N_{\text{AV}}$: We investigate the impact of the key hyperparameter $N_{\text{AV}}$ on CcGAN-AVAR-H’s performance (Figure 7), conducting this ablation study with auxiliary regularization disabled to isolate AV’s effect. The results demonstrate substantial performance degradation when either disabling AV entirely ($N_{\text{AV}}=0$) or employing an excessively wide vicinity (large $N_{\text{AV}}$ values), while moderate $N_{\text{AV}}$ values consistently yield acceptable results. The dashed green lines in Figure 7 correspond to the optimal values selected for our main experiments.
@@ B.2 Relation to ACGAN
CcGAN-AVAR differs fundamentally from ACGAN (Odena, Olah, and Shlens 2017) in three key aspects. First, while ACGAN uses an auxiliary classification branch for class-conditional generation, its discriminator does not directly receive class information. In contrast, CcGAN-AVAR integrates a regression branch for continuous label prediction and explicitly injects condition labels into the discriminator via label projection (Figure 3). Second, whereas ACGAN relies on a cross-entropy loss for classification, CcGAN-AVAR employs a regression loss—specifically, a custom $\gamma$-insensitive hinge loss (Eq. (6)) designed to improve robustness against label noise, rather than a standard mean squared error. Third, and most distinctively, CcGAN-AVAR implements a density ratio estimation branch that imposes an $f$-divergence penalty during generator training - a feature entirely absent in the standard ACGAN architecture.
@@ C.2 Hardware and Software Details
  • $64\times 64$ resolution: 24GB GPU memory
  • $128\times 128$ and $192\times 192$ resolutions: 48GB GPU memory (RTX 4090D-48G)
@@ Appendix D Extra Experimental Results
Example synthetic images generated by Dual-NDA and CcGAN-AVAR for the UTKFace and Steering Angle experiments ($\times128$ resolution) are presented in Figs. S.12 and S.13.
Figure S.12: Some example fake images from Dual-NDA and CcGAN-AVAR-H in the UTKFace ($\times128$) experiment.
Figure S.13: Some example fake images from Dual-NDA and CcGAN-AVAR-H in the Steering Angle ($\times128$) experiment.